Consider a competitive industry with two firms and the production level of y. One has cost function ?1 (?) = ? 2 , and other has cost function ?2 (?) = 2? 2 . Assume the output price is p. Find the industry supply function.
Answer #1

Answer

In order to maximize profit a perfect competitive firm produces that quantity at which P = MC where P = Output Price and MC = Marginal Cost.

FIRM 1 :

C1 = y2

MC1 = dC1/dy1 = 2y1 Thus P = MC1 => P = 2y1 => y1 = P/2 ----------------------Supply of firm 1

Here y1 = quantity of output of firm 1

FIRM 2 :

C2 = 2y2

MC2 = dC1/dy2 = 4y2 Thus P = MC2 => P = 4y2 => y2 = P/4 ----------------------Supply of firm 2

Here y2 = quantity of output of firm 2

Market or industry supply curve is the horizontal summation of individual supply curve.

Thus industry supply curve is given by :

y = y1 + y2 => P/2 + P/4

=> y = 0.75P ----------------------------Industry supply curve